OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x2 + 8x + 18

hero (hero):

Subtract 18 from both sides: y - 18 = x^2 + 8x Complete the square by adding 16 to both sides: y - 18 + 16 = x^2 + 8x + 16 y - 2 = x^2 + 8x + 16 Express x^2 + 8x + 16 in binomial square form: y - 2 = (x + 4)^2 Add 2 to both sides: y = (x + 4)^2 + 2

hero (hero):

Remember, the vertex form is y = a(x - h)^2 + k In this case a = 1, h = -4, k = 2

OpenStudy (anonymous):

Thanks! Could you help me out with this problem? I'm pretty sure the answer is the vertex of the function, which I got 20 for but I'm not totally sure. A projectile is thrown upward so that its distance above the ground after t seconds is given by the function h(t) = - 16t2 + 640t. After how many seconds does the projectile take to reach its maximum height?

hero (hero):

You can use the vertex formula to figure out the max height: \(t = -\dfrac{b}{2a}\) \(t = -\dfrac{640}{2(-16)}\) \(t = 20\) Next you'll need to insert t = 20 into the given equation: h(20) = - 16(20)^2 + 640(20)

hero (hero):

then simplify the expression on the right to solve for h(20)

OpenStudy (anonymous):

@Hero how would you self for h(20)

OpenStudy (anonymous):

*solve